1. Expected cash dividends are $2.50, the dividend yield is 7%, flotation costs are 5%, and the growth rate is 3%. Compute the cost of new common stock.

A. 10.37%

B. 8.89%

C. 9.48%

D. 9.34%

2. Debreu Beverages has an optimal capital structure that is 50% common equity, 40% debt, and 10% preferred stock. Debreu's pretax cost of equity is 12%. Its pretax cost of preferred stock is 7%, and its pretax cost of debt is also 7%. If the corporate tax rate is 35%, what is the weighed average cost of capital?

A. 8.52%

B. 8.67%

C. 9.50%

D. 7.68%
Answer #1

1)

Current price =2.5/7% =35.71

Cost of new stock =Expected dividend/current price+growth

Cost of new stock =(2.5/(35.71*95%))+3%

Cost of new stock =10.37%

2)

After tax cost of debt =7%*(1-35%) =4.55%

WACC =(50%*12%)+(40%*4.55%)+(10%*7%)

WACC =8.52%